# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 5.1. What transfers from crowdsourcing
Crowdsourcing studies currently support a wider range of tasks for creative work, under the verifiable creativity and creativity objectives, with a wider range of approaches. We expect these approaches to transfer to chaining and bring myriad benefits.
@@
Fewer studies include creative tasks in chaining (15 papers or 38.5%) than in crowdsourcing (51 papers or 75%). Future work could attempt creative tasks that are under-explored in chaining, such as: prototyping (Lee et al., 2017; Valentine et al., 2017; Kulkarni et al., 2014), categorizing (Huffaker et al., 2020; Bragg et al., 2013; Chang et al., 2017), topic concept maps (Liu et al., 2018; Hahn et al., 2016; Luther et al., 2015), project development (Mahyar et al., 2018), or more specific tasks like itinerary creation (Zhang et al., 2012) and conceptual blends (Chilton et al., 2019).
@@
Crowdsourcing work also uses a broader range of communicative and dynamic architectures to support creative work. Giving workers roles during communication can assist both crowdsourcing workflows (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014) and LLM chains (Wu et al., 2023a). In crowdsourcing work, some studies go beyond giving a role in instructions and instead create free-form role-based (Nebeling et al., 2016; Valentine et al., 2017; Zhang et al., 2023b) hierarchical (Agapie et al., 2015; Retelny et al., 2014) and state-sharing (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013) communicative architectures.
@@
5.2. What does not transfer from crowdsourcing
Although strategic aspects of workflow task decomposition transfer, the differences between crowdworkers and LLMs lead to different optimal tactical design choices in the selection and implementation of architectures and subtasks. Examples of tactical design decisions include determining the number of optimal parallel generations, performing a series of single verification subtasks instead of one multi-faceted subtask, and the text used for worker instructions in a subtask. Compared to crowdworkers, LLMs can process and generate fluent text more quickly per subtask. They best respond to different, often more constrained, task instructions and require more direction to generate and verify with diverse responses. We highlight several instances in which these differences affect optimal tactical design choices and studies that could guide future work towards understanding how LLMs respond to different tactical design choices.
5.2.1. Sufficient context must not overcomplicate subtasks.
Providing global context to subtasks is widely agreed to be a critical element of subtask design and output coherence, but it can increase instruction complexity (Wu et al., 2022a; Salehi et al., 2017; Ambati et al., 2012; Luther et al., 2015). However, initial findings from the literature survey and case studies suggest that LLMs can respond poorly to long, complicated instructions (Du et al., 2023; Parameswaran et al., 2023; Madaan et al., 2023b; Wu et al., 2023c). This context limitation should be re-evaluated over time as LLM context windows grow.
For example, we wanted to provide Mechanical Novel the entire short story thus far as context, as occurs in Crowdsourced Mechanical Novel. To generate diverse suggestions for improvement we gave the LLM previous suggestions as context for what it should avoid. However, adding this requirement to the overall prompt anecdotally increased the output’s chance of anchoring to prior suggestions. Therefore, we made a separate evaluate subtask that determined if the new suggestion matched previous suggestions out of context from the rest of the story.
We recommend designers take into consideration that LLMs may not adequately process all the nuances of complicated instructions. Simplifying the subtask size in turn simplifies instructions.
5.2.2. LLMs need different, often simpler, subtasks
Throughout our case studies, we often needed to break crowdsourcing subtasks into further subtasks. Multiple factors necessitated these changes: natural language is difficult to structure (e.g., thus, we query for one item category at a time rather than eight in Cascade), LLMs have difficulty balancing multiple requirements in a subtask (Wu et al., 2023c; Madaan et al., 2023b) (e.g., in response, we split Soylent’s “Fix” step into separate edit, merge, and delete queries), and LLMs do not understand length-related constraints (e.g., to address this, we add a hard-coded verification step after a generate subtask in Mechanical Novel to check for length).
In our case studies, LLMs performed better on highly specified subtasks for chains rather than fewer, more complicated subtasks.
5.2.3. Sourcing diverse responses requires more overhead but may offer more control.
Crowdsourcing workflows get diverse responses by querying the crowdsourcing platform multiple times and receiving the suggestion of a different worker each time. LLMs have the capability to output diverse responses, as their training data includes information produced by a wide variety of people. Although LLMs give stochastic responses, our initial case study experiences suggest that querying an LLM multiple times may output responses with less diversity than querying multiple crowdworkers. Methods for sourcing diverse responses include prompt diversity, verification checks to judge uniqueness, temperature adjustments, and using a variety of models. Parallel and iterative architectures source different degrees of diversity and quality of responses in crowdsourcing (Little et al., 2010a). Although these takeaways may not transfer directly, redundant architecture design choices will also impact LLM response diversity.
Although sourcing diverse responses requires more specific prompt and workflow design with LLMs, initial results from role-based prompts suggest an opportunity for finer-grained control of LLMs (Wu et al., 2023a; Zhang et al., 2023b; Wang et al., 2023a). Critical challenges remain around what tactical design decisions can generate usefully diverse responses with LLMs. Future studies can measure the impact of these design decisions to more effectively harness the response diversity strategy.
To source diverse responses, we recommend further studies on how prompt diversity, model choices, role-based prompts, and architectures can source diverse responses from LLMs.
5.2.4. Simple verification methods are often insufficient.
Providing verification in crowdsourcing workflows often involves requesting a different worker to verify if errors or desirable qualities exist in outputs. Some studies show that prompting a model to reflect can help achieve higher accuracy outputs (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023). However, others caution against the same model verifying its own outputs, as LLMs exhibit high confidence for incorrect prior outputs (Liang et al., 2023; Du et al., 2023). We found mixed results through our case studies on the effectiveness of a model verifying its own output. These mixed results from the literature and our case studies imply that more robust verification methods could be especially useful when working with LLMs. Some approaches for more robust verification include: verifying with a different model (Liang et al., 2023), debate-based structures (Liang et al., 2023; Du et al., 2023), structural verification such as similarity across parallel streams (Du et al., 2023; Liem et al., 2011), sets of specific evaluate subtasks, and hard-coded verification.
Compared to crowdsourcing, we recommend more robust methods for verifying quality, such as model variety and debate-based communicative architectures.
5.2.5. LLM chains need new design studies.
@@ 5.2.2. LLMs need different, often simpler, subtasks
These observations are some of many differences in design requirements. More systematic studies on optimal chain design would be valuable. Several studies already approach questions around model choice (Madaan et al., 2023a; Du et al., 2023), incorporating roles into prompts (Wu et al., 2023a; Park et al., 2022), and using self-reflection or debate to validate outputs (Du et al., 2023; Liang et al., 2023). Although the outcomes may be different, systematic crowdsourcing studies could provide inspiration for some of these explorations. Prior studies investigate the effect of context timing, architecture, and content on task output (Luther et al., 2015; Salehi et al., 2017). Others measure trade-offs of iterative and parallel workflows (e.g., effect on latency, quality averages and variance, risk of cascading errors) in various task contexts, as well as good practices for each (e.g., appropriate number of branches or iterations and what context to share amongst branches) (Kittur et al., 2011; Chang et al., 2017; Luther et al., 2015; Huffaker et al., 2020; Little et al., 2010a; Goto et al., 2016; André et al., 2014).
@@
By specifying a strategy of interest, the design space can orient studies for understanding optimal tactical design decisions. For example, studying debate designs could look at debate to establish flexibility for tasks with many edge cases, debate to resolve disagreements, or debate to encourage a diversity of creative ideas. Evaluating the use of tactic designs within the context of specific strategies will lead to more generalizable findings than evaluating solely on task accuracy.
@@
We recommend systematic tests for tactical design decisions informed by prior crowdsourcing work and the design space.
5.3. New opportunities for LLM chain design
@@
Since LLMs are cheaper and faster to query (Terwiesch and Ulrich, [n. d.]), there are more opportunities to find optimal designs through systematic testing. While the lower barrier to entry may allow LLM chain designers to iterate more quickly, it doesn’t guarantee that they focus their attention on the right things while iterating. Since LLMs display unpredictable outputs in response to low-level changes, designers may end up fixating on minor tweaks like prompt rewordings rather than exploring the design space of strategies and tactics. To avoid potential design fixation that tunnels into technical detail rather than exploring higher-level design directions, LLM chain designers would benefit from tools that scaffold their design and iteration process. Running an LLM chain is cheaper and faster than running a crowdsourcing workflow, but the design process is unstructured. There is a need for design support.
@@
As there are many overlapping tactics among workflows, a repository of “battle-tested” workflow operators could allow sharing of quality parts for reuse in new workflows. Such a platform could include operator-level test suites to assist with validation—and could draw on crowdsourcing to create such test suites—that would monitor operator output quality in new content areas or with newly released LLMs. Additionally, corresponding workflow design tools could suggest different options to achieve a goal by collecting multiple approaches for sub-parts of the workflow.
Other work engineers more optimal LLM prompts by iteratively generating and scoring test prompt candidates using mathematical or natural language optimization functions (Zhou et al., 2022a; Yang et al., 2023). Future work could adapt these approaches within the context of chaining, for example to optimize for diverse answers.
Comparing many workflow designs can be difficult because of the requirement that the output of a previous step match the input of the next step. Therefore, removing or adding steps can be a non-trival process. More formalized input and output indicators, potentially with matching prompts, could facilitate easier chain development. This approximates the approach of Flash Teams which documents the inputs and outputs of steps (Retelny et al., 2014).
@@
5.4. Related work on LLMs and crowdsourcing
Concurrent to this paper, a few other groups investigate the intersection of crowdsourcing and LLMs. One paper explores the impact of incorporating LLMs into crowdsourcing workflows on requester and crowdworker stakeholders (ALLEN et al., 2023). Another implements three strategies from crowdsourcing using LLMs for data processing tasks (Parameswaran et al., 2023). Closest to our work, the final paper has students implement six crowdsourcing workflows. The paper discusses the effectiveness of these workflows and the impact of the difference between LLM and crowdworker abilities on recommendations for chaining (Wu et al., 2023c). We build on this approach of providing recommendations based on case study implementations. We widen the scope of our recommendations by first integrating a broad survey of LLM chaining and crowdsourcing workflow literature into a design space and then using case studies to explore open questions in the design space. This broader investigation enables us to build a design space and informs more comprehensive recommendations for navigating design decisions and for orienting future work. These recommendations span from higher-level guidance, such as adhering to the quality-cost tradeoff, to lower-level suggestions, such as pursuing architecture-based methods for quality verification.
@@ 6. Limitations
Broader impacts. We focus on the ability of workflows to serve a beneficial role in supporting various tasks. However, there are potential risks in using LLMs and crowdworkers as support tools that we do not explicitly consider. Several papers in the design space cite concerns around the loss of human jobs and unintended consequences of using LLMs, especially for creative tasks (Park et al., 2022; Mirowski et al., 2023; Wu et al., 2023a). Crowdsourcing work similarly addressed concerns around replacing expert workers and labor rights for crowdworkers (Kittur et al., 2013). Broader discussions about privacy (Wu et al., 2023a; Peris et al., 2023), ownership (Kim et al., 2023b; Mirowski et al., 2023), and perceived sense of agency we also leave to future work. Finally, although workflows can improve output quality, results will still be imperfect, so errors from implemented chains could still have harmful impacts (Eggmann et al., 2023; Greshake et al., 2023).
@@ 7. Conclusion
We argue that crowdsourcing workflows can be a fruitful source of inspiration for designing LLM chains. We contribute a design space for LLM chain designers that incorporates lessons from the crowdsourcing workflow literature. This design space both encourages an objectives-based approach to chain design and opens questions about what crowdsourcing decomposition strategies can transfer to LLMs. Through three case studies, we address this knowledge gap about the design space by building an understanding for how crowdsourcing workflows need to be adapted for LLM chains. Finally, we synthesize our findings from the literature and our case study to propose that the categories of the design space transfer, but that different tactical design decisions can be needed to support LLM chains. We find LLM chains need additional support in subtask design and verification, but also that their comparatively low cost and high speed can allow for faster iteration of designs than in crowdsourcing. Through these findings, we guide future work towards more approaches to user revision and creative tasks, systematic discovery of good practices for tactic design, and artifacts to support the chain design process.
@@ Acknowledgements.
We thank Michael Bernstein and Joy Kim for providing outputs from the original Soylent and Mechanical Novel papers. We also thank the Allen Institute for Artificial Intelligence for access to Amazon Mechanical Turk resources. We thank OpenAI for offering research credits to the GPT-3 API. This research was supported in part by NSF award IIS-1901386, ONR grant N00014-21-1-2707, and the Allen Institute for Artificial Intelligence (AI2).References
@@
- Wang et al. (2023b) Sitong Wang, Savvas Petridis, Taeahn Kwon, Xiaojuan Ma, and Lydia B Chilton. 2023b. PopBlends: Strategies for conceptual blending with large language models. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–19.
- Wang et al. (2023a) Zekun Moore Wang, Zhongyuan Peng, Haoran Que, Jiaheng Liu, Wangchunshu Zhou, Yuhan Wu, Hongcheng Guo, Ruitong Gan, Zehao Ni, Man Zhang, et al. 2023a. Rolellm: Benchmarking, eliciting, and enhancing role-playing abilities of large language models. arXiv preprint arXiv:2310.00746 (2023).

### v3

@@ 8.1. Creative tasks with dynamic and communicative architectures
Crowdsourcing studies currently support a wider range of tasks for creative work with a wider range of approaches than chaining studies. We expect these approaches to transfer to chaining and bring myriad benefits.
@@
Future work could attempt creative tasks that are under-explored in chaining, such as: prototyping (Lee et al., 2017; Valentine et al., 2017; Kulkarni et al., 2014), categorizing (Huffaker et al., 2020; Bragg et al., 2013; Chang et al., 2017), topic concept maps (Liu et al., 2018; Hahn et al., 2016; Luther et al., 2015), project development (Mahyar et al., 2018), or more specific tasks like itinerary creation (Zhang et al., 2012), researching a purchase (Kittur et al., 2011), and conceptual blends (Chilton et al., 2019).
@@
Crowdsourcing work also uses a broader range of communicative and dynamic architectures to support creative work. Giving workers roles during communication can assist both crowdsourcing workflows (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014) and LLM chains (Wu et al., 2023a). In crowdsourcing work, some studies go beyond giving a role in instructions and instead create free-form role-based (Nebeling et al., 2016; Valentine et al., 2017; Zhang et al., 2023b), hierarchical (Agapie et al., 2015; Retelny et al., 2014), and state-sharing (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013) communicative architectures.
@@
8.2. Integrating the user
Although many chaining papers discuss involving humans into the chain hypothetically, few implement this user oversight (Stretcu et al., 2023; Toubal et al., 2024). Our case studies showcase one example of interactions through directly manipulable outputs, but the varied ways of incorporating the user from crowdsourcing workflows could continue to inspire similar approaches in chaining.
We recommend incorporating more approaches to user guidance from crowdsourcing, such as explorable interfaces (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019), worker-initiated revision (Salehi et al., 2017), and guided prototyping (Nebeling et al., 2016; Lasecki et al., 2015). We also recommend research into the the opportunities and limitations of user guidance, especially for creative tasks, as initial findings conflict on if using generative AI tools leads to divergent thinking (chi, 2023) or design fixation (Wadinambiarachchi et al., 2024).
8.3. Studying chains more systematically
@@
More systematic studies on optimal chain design would be valuable. Several studies already approach questions around model choice (Madaan et al., 2023a; Du et al., 2023), incorporating roles into prompts (Wu et al., 2023a; Park et al., 2022), and using self-reflection or debate to validate outputs (Du et al., 2023; Liang et al., 2023). Although the outcomes may be different, systematic crowdsourcing studies could provide inspiration for some of these explorations. Prior crowdsourcing studies investigate the effect of context timing, architecture, and content on task output (Luther et al., 2015; Salehi et al., 2017). Others measure tradeoffs of iterative and parallel workflows (e.g., effect on latency, quality averages and variance, risk of cascading errors) in various task contexts, as well as good practices for each (e.g., appropriate number of branches or iterations and what context to share amongst branches) (Kittur et al., 2011; Chang et al., 2017; Luther et al., 2015; Huffaker et al., 2020; Little et al., 2010a; Goto et al., 2016; André et al., 2014).
@@
By specifying a strategy of interest, the design space can orient studies for understanding optimal tactical design decisions. For example, studying debate designs could look at debate to establish adaptability for tasks with many edge cases, debate to validate outputs, or debate to encourage a diversity of creative ideas. Evaluating the use of tactic designs within the context of specific strategies will lead to more generalizable findings than evaluating solely on task performance.
@@
We recommend systematic tests for tactical design decisions informed by prior crowdsourcing work, objectives, and strategies.
8.4. LLM workflow design processes are unstructured and design support is needed
@@
Since LLMs are cheaper and faster to query (Terwiesch and Ulrich, [n. d.]), there are more opportunities to find optimal designs through systematic testing. While the lower barrier to entry may allow LLM chain designers to iterate more quickly, it doesn’t guarantee that they focus their attention on the right things while iterating. Since LLMs display unpredictable outputs in response to low-level changes, designers may end up fixating on minor tweaks like prompt rewordings rather than exploring the design space of strategies and tactics. To avoid potential design fixation that tunnels into technical detail rather than exploring higher-level design directions, LLM chain designers would benefit from tools that scaffold their design and iteration process. Although running an LLM chain is cheaper and faster than running a crowdsourcing workflow, the design process is unstructured. There is a need for design support.
@@
As there are many overlapping tactics among workflows, the flexibility and reusability of modules can support the development of future workflows (Josifoski et al., 2023). A repository of “battle-tested” workflow operators could allow sharing of quality parts for reuse in new workflows. Such a platform could include operator-level test suites to assist with validation that would monitor operator output quality in new content areas or with newly released LLMs. The platform design could draw on crowdsourcing to create such test suites. Additionally, corresponding workflow design tools could suggest different options to achieve a goal by collecting multiple approaches for sub-parts of the workflow.
Other work engineers more optimal LLM prompts by iteratively generating and scoring test prompt candidates using mathematical or natural language optimization functions (Zhou et al., 2022a; Yang et al., 2023). Future work could adapt these approaches within the context of chaining, for example to optimize for diverse responses.
Comparing many workflow designs can be difficult because of the requirement that the output of a previous step match the input of the next step. Therefore, removing or adding steps can be a non-trivial process. More formalized input and output indicators, potentially with matching prompts, could facilitate easier chain development. This approximates the approach of Flash Teams which documents the inputs and outputs of steps (Retelny et al., 2014).
@@ 9. Limitations
Broader impacts. We focus on the ability of workflows to serve a beneficial role in supporting various tasks. However, there are potential risks in using LLMs and crowdworkers as support tools that we do not explicitly consider. Several papers in the design space cite concerns around a loss of jobs and unintended consequences of using LLMs, especially for creative tasks (Park et al., 2022; Mirowski et al., 2023; Wu et al., 2023a). Crowdsourcing work similarly raised concerns around replacing expert workers and around labor rights for crowdworkers (Kittur et al., 2013). Broader discussions about privacy (Wu et al., 2023a; Peris et al., 2023), ownership (Kim et al., 2023b; Mirowski et al., 2023), and perceived sense of agency we also leave to future work. Finally, although workflows can improve output quality, results will still be imperfect, so errors from implemented chains could still have harmful impacts (Eggmann et al., 2023; Greshake et al., 2023).
@@ 10. Conclusion
By revisiting the established field of crowdsourcing workflows, we find inspiration and direction for the nascent area of LLM chaining. To facilitate adaptation, we contribute a design space for LLM chain designers that incorporates findings from the crowdsourcing workflow and LLM chaining literatures. This design space covers the objectives supported by and the tactical building blocks of prior workflows and chains. We also surface strategies employed by designers to use tactics to support objectives. Through three case studies, we build an understanding for how crowdsourcing workflows can be adapted for LLM chains.
Finally, we synthesize our findings to discuss tactical design decisions for LLM chains through the context of their comparison with crowdsourcing workflows. We find LLM chains need additional support in subtask design and validation due in part to differences between LLM stochasticity and human diversity. On the other hand, the comparatively low cost and latency of LLM calls allows for faster iteration and for workflow techniques that were studied but never practical with human crowds. We hope these findings guide future work towards more approaches to user guidance and creative tasks, systematic discovery of good practices for tactic design, and artifacts to support the chain design process.
LLM chains present remarkable flexibility, which comes along with both promising capabilities but also substantial complexity. Our work helps LLM chain designers to navigate this vast design space by building on the established field of crowdsourcing.
@@ Acknowledgements.
We thank Michael Bernstein and Joy Kim for providing outputs from the original Soylent and Mechanical Novel papers. We also thank the Allen Institute for Artificial Intelligence for access to Amazon Mechanical Turk resources. We thank OpenAI for offering research credits to the GPT-3.5 API. This research was supported in part by NSF award IIS-1901386, ONR grant N00014-21-1-2707, and the Allen Institute for Artificial Intelligence (AI2).References
@@
- chi (2023) 2023. DIS ’23: Proceedings of the 2023 ACM Designing Interactive Systems Conference (, Pittsburgh, PA, USA,). Association for Computing Machinery, New York, NY, USA.
@@
- Alexander (1977) Christopher Alexander. 1977. A Pattern Language: Towns, Buildings, Construction. Oxford University Press.
@@
- Bender et al. (2021) Emily M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big?. In Proceedings of the 2021 ACM conference on fairness, accountability, and transparency. 610–623.
@@
- Difallah et al. (2018) Djellel Difallah, Elena Filatova, and Panos Ipeirotis. 2018. Demographics and dynamics of mechanical turk workers. In Proceedings of the eleventh ACM international conference on web search and data mining. 135–143.
@@
- Gamma et al. (1994) Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. 1994. Design Patterns Elements of Reusable Object-Oriented Software. Addison-Wesley.
@@
- Gordon et al. (2022) Mitchell L. Gordon, Michelle S. Lam, Joon Sung Park, Kayur Patel, Jeff Hancock, Tatsunori Hashimoto, and Michael S. Bernstein. 2022. Jury Learning: Integrating Dissenting Voices into Machine Learning Models. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (, New Orleans, LA, USA,) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 115, 19 pages. https://doi.org/10.1145/3491102.3502004
@@
- Horvitz (1999) Eric Horvitz. 1999. Principles of mixed-initiative user interfaces. In Proceedings of the SIGCHI conference on Human Factors in Computing Systems. 159–166.
- Hsieh et al. (2023) Cheng-Yu Hsieh, Si-An Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023. Tool documentation enables zero-shot tool-usage with large language models. arXiv preprint arXiv:2308.00675 (2023).
@@
- Josifoski et al. (2023) Martin Josifoski, Lars Klein, Maxime Peyrard, Yifei Li, Saibo Geng, Julian Paul Schnitzler, Yuxing Yao, Jiheng Wei, Debjit Paul, and Robert West. 2023. Flows: Building blocks of reasoning and collaborating ai. arXiv preprint arXiv:2308.01285 (2023).
@@
- Li (2024) Jiyi Li. 2024. A Comparative Study on Annotation Quality of Crowdsourcing and LLM via Label Aggregation. CoRR abs/2401.09760 (2024). https://doi.org/10.48550/ARXIV.2401.09760 arXiv:2401.09760
@@
- Liu et al. (2023) Ryan Liu, Howard Yen, Raja Marjieh, Thomas L Griffiths, and Ranjay Krishna. 2023. Improving interpersonal communication by simulating audiences with language models. arXiv preprint arXiv:2311.00687 (2023).
@@
- Sorensen et al. (2024) Taylor Sorensen, Jared Moore, Jillian Fisher, Mitchell Gordon, Niloofar Mireshghallah, Christopher Michael Rytting, Andre Ye, Liwei Jiang, Ximing Lu, Nouha Dziri, et al. 2024. A Roadmap to Pluralistic Alignment. arXiv preprint arXiv:2402.05070 (2024).
@@
- Toubal et al. (2024) Imad Eddine Toubal, Aditya Avinash, Neil Gordon Alldrin, Jan Dlabal, Wenlei Zhou, Enming Luo, Otilia Stretcu, Hao Xiong, Chun-Ta Lu, Howard Zhou, Ranjay Krishna, Ariel Fuxman, and Tom Duerig. 2024. Modeling Collaborator: Enabling Subjective Vision Classification With Minimal Human Effort via LLM Tool-Use. Proceedings of the IEEE/CVF International Conference on Computer Vision (2024).
@@
- Wadinambiarachchi et al. (2024) Samangi Wadinambiarachchi, Ryan M Kelly, Saumya Pareek, Qiushi Zhou, and Eduardo Velloso. 2024. The Effects of Generative AI on Design Fixation and Divergent Thinking. arXiv preprint arXiv:2403.11164 (2024).
- Wang et al. (2024) Angelina Wang, Jamie Morgenstern, and John P Dickerson. 2024. Large language models cannot replace human participants because they cannot portray identity groups. arXiv preprint arXiv:2402.01908 (2024).
@@
- Wang et al. (2023) Sitong Wang, Savvas Petridis, Taeahn Kwon, Xiaojuan Ma, and Lydia B Chilton. 2023. PopBlends: Strategies for conceptual blending with large language models. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–19.
@@
- Zhang et al. (2024) Shaokun Zhang, Jieyu Zhang, Jiale Liu, Linxin Song, Chi Wang, Ranjay Krishna, and Qingyun Wu. 2024. Training Language Model Agents without Modifying Language Models. International Conference on Machine Learning (2024).